# Ratio Covers of Convex Sets and Optimal Mixture Density Estimation

### v1

@@ D.5 Proof of Proposition˜4
We fix $\rho\in(0,1]$ and let $\mathcal{A}_{\rho}$ be the minimal $\rho$-net of $\Delta_{M-1}$ with respect to the $\|\cdot\|_{1}$ distance. Define the finite subclass $\mathcal{W}_{\rho}=\{w_{\alpha}:\alpha\in\mathcal{A}_{\rho}\}\subseteq\mathcal{W}$. Let $p$ be any element in $\mathcal{W}$ and $p_{\rho}$ be the $\rho$-close element to it in $\mathcal{W}_{\rho}$ with respect to the net $\mathcal{A}_{\rho}$. By Lemma˜12 and (30) we have
$Z(p)-Z(p_{\rho})\leqslant 4nM\rho+\frac{\log(2nM)}{(\log(2nM)+1)}nM\rho\leqslant 5nM\rho,$
@@
${\mathbf{E}}_{\varepsilon}\exp\Big(\lambda\sup_{p\in\mathcal{W}}Z(p)\Big)$ $\leqslant\exp\Big(5\lambda nM\rho\Big)\,{\mathbf{E}}_{\varepsilon}\exp\Big(\lambda\max_{q\in\mathcal{W}_{\rho}}Z(q)\Big)$
$\leqslant|\mathcal{W}_{\rho}|\exp\Big(5\lambda nM\rho\Big)\max\limits_{p\in\mathcal{W}_{\rho}}{\mathbf{E}}_{\varepsilon}\exp\Big(\lambda Z(p)\Big).$
Using ${\mathbf{E}}\exp(\lambda\varepsilon_{i})\leqslant\exp(\lambda^{2}/2)$ we have for any $\lambda\in\left[0,\frac{n}{16(\log(2nM)+1)}\right]$ and $p\in\mathcal{W}$,
${\mathbf{E}}_{\varepsilon}\exp\Big(\lambda Z(p)\Big)$ $={\mathbf{E}}_{\varepsilon}\exp\left(\frac{\lambda}{n}\sum\limits_{i=1}^{n}\left(4\varepsilon_{i}\log\left(\frac{p^{\star}_{\mathcal{W}}(X_{i})}{p(X_{i})}\right)-\frac{1}{2(\log(2nM)+1)}\log^{2}\left(\frac{p^{\star}_{\mathcal{W}}(X_{i})}{p(X_{i})}\right)\right)\right)$
$\leqslant\exp\left(\sum\limits_{i=1}^{n}\left(\frac{8\lambda^{2}}{n^{2}}\log^{2}\left(\frac{p^{\star}_{\mathcal{W}}(X_{i})}{p(X_{i})}\right)-\frac{\lambda}{2n(\log(2nM)+1)}\log^{2}\left(\frac{p^{\star}_{\mathcal{W}}(X_{i})}{p(X_{i})}\right)\right)\right)$
@@
${\mathbf{E}}\exp\Big(\lambda\big(L(\widehat{p})-L(p^{\star}_{\mathcal{W}})\big)\Big)\leqslant|\mathcal{W}_{\rho}|\exp\Big(5\lambda nM\rho\Big)\leqslant\exp\left(M\log(3/\rho)+5\lambda nM\rho\right),$ (36)
where we used the standard volumetric argument (see [58, Lemma 5.7]) to bound $|\mathcal{W_{\rho}}|$ for $\rho\leqslant 1$. This immediately implies that, with probability at least $1-\delta$,
$L(\widehat{p})-L(p^{\star}_{\mathcal{W}})\leqslant 5nM\rho+\frac{1}{\lambda}\Big(M\log(3/\rho)+\log(1/\delta)\Big).$ (37)
Taking $\lambda=\frac{n}{16(\log(2nM)+1)}$ and $\rho=\frac{1}{3n^{2}M}\leqslant 1$, we obtain, using $L(\widehat{p})-L(p^{\star}_{\mathcal{W}})=\operatorname{KL}(p^{\star},\widehat{p})-\operatorname{KL}(p^{\star},p^{\star}_{\mathcal{W}})$,
$\operatorname{KL}(p^{\star},\widehat{p})-\operatorname{KL}(p^{\star},p^{\star}_{\mathcal{W}})\leqslant\frac{5}{3n}+\frac{16(\log(2nM)+1)(M\log(9n^{2}M)+\log(1/\delta))}{n}.$ (38)
@@
$\operatorname{KL}(p^{\star},p^{\star}_{\mathcal{W}})\leqslant\inf_{p\in\mathop{\mathrm{conv}}(\mathcal{P})}\operatorname{KL}(p^{\star},p)+\frac{2}{n}.$ (39)

### v2

@@ D.5 Proof of Proposition˜4
We fix $\rho\in(0,1]$ and let $\mathcal{A}_{\rho}$ be the minimal $\rho$-net of $\Delta_{M-1}$ with respect to the $\|\cdot\|_{1}$ distance. Define the finite subclass $\mathcal{Q}_{\rho}=\{w_{\alpha}:\alpha\in\mathcal{A}_{\rho}\}\subseteq\mathcal{Q}$. Let $p$ be any element in $\mathcal{Q}$ and $p_{\rho}$ be the $\rho$-close element to it in $\mathcal{Q}_{\rho}$ with respect to the net $\mathcal{A}_{\rho}$. By Lemma˜13 and (35) we have
$Z(p)-Z(p_{\rho})\leqslant 4nM\rho+\frac{\log(2nM)}{\log(2nM)+1}nM\rho\leqslant 5nM\rho,$
@@
${\mathbf{E}}_{\varepsilon}\exp\Big(\lambda\sup_{p\in\mathcal{Q}}Z(p)\Big)$ $\leqslant\exp\Big(5\lambda nM\rho\Big)\,{\mathbf{E}}_{\varepsilon}\exp\Big(\lambda\max_{q\in\mathcal{Q}_{\rho}}Z(q)\Big)$
$\leqslant|\mathcal{Q}_{\rho}|\exp\Big(5\lambda nM\rho\Big)\max\limits_{p\in\mathcal{Q}_{\rho}}{\mathbf{E}}_{\varepsilon}\exp\Big(\lambda Z(p)\Big).$
Using ${\mathbf{E}}\exp(\lambda\varepsilon_{i})\leqslant\exp(\lambda^{2}/2)$ we have for any $\lambda\in\left[0,\frac{n}{16(\log(2nM)+1)}\right]$ and $p\in\mathcal{Q}$,
${\mathbf{E}}_{\varepsilon}\exp\Big(\lambda Z(p)\Big)$ $={\mathbf{E}}_{\varepsilon}\exp\left(\frac{\lambda}{n}\sum\limits_{i=1}^{n}\left(4\varepsilon_{i}\log\left(\frac{p^{\star}_{\mathcal{Q}}(X_{i})}{p(X_{i})}\right)-\frac{1}{2(\log(2nM)+1)}\log^{2}\left(\frac{p^{\star}_{\mathcal{Q}}(X_{i})}{p(X_{i})}\right)\right)\right)$
$\leqslant\exp\left(\sum\limits_{i=1}^{n}\left(\frac{8\lambda^{2}}{n^{2}}\log^{2}\left(\frac{p^{\star}_{\mathcal{Q}}(X_{i})}{p(X_{i})}\right)-\frac{\lambda}{2n(\log(2nM)+1)}\log^{2}\left(\frac{p^{\star}_{\mathcal{Q}}(X_{i})}{p(X_{i})}\right)\right)\right)$
@@
${\mathbf{E}}\exp\Big(\lambda\big(L(\widehat{p})-L(p^{\star}_{\mathcal{Q}})\big)\Big)\leqslant|\mathcal{Q}_{\rho}|\exp\Big(5\lambda nM\rho\Big)\leqslant\exp\left(M\log(3/\rho)+5\lambda nM\rho\right),$ (41)
where we used the standard volumetric argument (see [61, Lemma 5.7]) to bound $|\mathcal{Q}_{\rho}|$ for $\rho\leqslant 1$. This immediately implies that, with probability at least $1-\delta$,
$L(\widehat{p})-L(p^{\star}_{\mathcal{Q}})\leqslant 5nM\rho+\frac{1}{\lambda}\Big(M\log(3/\rho)+\log(1/\delta)\Big).$ (42)
Taking $\lambda=\frac{n}{16(\log(2nM)+1)}$ and $\rho=\frac{1}{3n^{2}M}\leqslant 1$, we obtain, using $L(\widehat{p})-L(p^{\star}_{\mathcal{Q}})=\operatorname{KL}(p^{\star},\widehat{p})-\operatorname{KL}(p^{\star},p^{\star}_{\mathcal{Q}})$,
$\operatorname{KL}(p^{\star},\widehat{p})-\operatorname{KL}(p^{\star},p^{\star}_{\mathcal{Q}})\leqslant\frac{5}{3n}+\frac{16(\log(2nM)+1)(M\log(9n^{2}M)+\log(1/\delta))}{n}.$ (43)
@@
$\operatorname{KL}(p^{\star},p^{\star}_{\mathcal{Q}})\leqslant\inf_{p\in\mathop{\mathrm{conv}}(\mathcal{P})}\operatorname{KL}(p^{\star},p)+\frac{2}{n}.$ (44)